# Private Evolution Converges

### v1

@@ Abstract
Private Evolution (PE) is a promising training-free method for differentially private (DP) synthetic data generation. While it achieves strong performance in some domains (e.g., images and text), its behavior in others (e.g., tabular data) is less consistent. To date, the only theoretical analysis of the convergence of PE depends on unrealistic assumptions about both the algorithm’s behavior and the structure of the sensitive dataset. In this work, we develop a new theoretical framework to explain PE’s practical behavior and identify sufficient conditions for its convergence. For $d$-dimensional sensitive datasets with $n$ data points from a bounded domain, we prove that PE produces an $(\varepsilon,\delta)$-DP synthetic dataset with expected 1-Wasserstein distance $\tilde{O}(d(n\varepsilon)^{-1/d})$ from the original, establishing worst-case convergence of the algorithm as $n\to\infty$. Our analysis extends to general Banach spaces as well. We also connect PE to the Private Signed Measure Mechanism, a method for DP synthetic data generation that has thus far not seen much practical adoption. We demonstrate the practical relevance of our theoretical findings in simulations.
@@ 1 Introduction
Many modern machine learning applications rely on user data, making data privacy protection a central concern. In this regard, differential privacy (DP) [16, 18] has become the primary standard to safeguard sensitive information of individuals. A wide range of problems including regression [37, 11], deep learning [1], and stochastic optimization [6, 2] can be performed in a DP manner. While it is possible to design a different DP algorithm for each specific task, a reasonable alternative is to generate a differentially private synthetic dataset that preserves many of the statistical properties of the sensitive dataset. By the post-processing property of DP, the DP synthetic data can then be fed into any existing non-DP algorithms, avoiding the need to modify existing ML pipelines, and can be safely shared with third parties (e.g. for reproducibility purposes).
Recently, [31] introduced Private Evolution (PE), a promising new framework for DP synthetic data generation that relies on public, pretrained data generators [44, 30, 26, 38, 46, 27]. PE is currently competitive with — and sometimes improves on — state-of-the-art models in terms of Fréchet inception distance (FID) and downstream task performance in settings such as images and text [31, 44, 30, 26]. In addition, PE is training-free, whereas current state-of-the-art approaches typically train (or finetune) a generative model on the sensitive dataset using DP-SGD [45, 33, 13, 10]. However, in some settings, including tabular data [38], and image data with mismatched distributions [20], PE has achieved limited success. To better understand when PE works, it is crucial to improve our theoretical understanding of the algorithm.
@@
In [31], the authors study a theoretically tractable variant of PE, analyzing its convergence in terms of the Wasserstein distance, with the aim of understanding its behavior and justifying its empirical success. However, this theoretically tractable version of PE differs in many significant ways from the algorithm that is used in experiments; further, the convergence analysis in [31] makes the unrealistic assumption that every point in $S$ is repeated many times. We start by showing that the proof technique in [31] is inherently limited by this multiplicity assumption; if we remove the assumption, then PE can only be $\varepsilon$-DP with a parameter $\varepsilon$ that scales as $d\log(1/\eta)$, where $d$ is the ambient dimension of the data and $\eta$ is the final accuracy of the algorithm in terms of the Wasserstein distance. This is impractical in even moderate dimensions, so new ideas are required to theoretically analyze utility.
In this paper, we address the limitations in prior analysis of PE by providing a convergence analysis that does not require strong assumptions about the nature of the dataset and more closely matches the PE algorithm used in practice. We formally prove worst-case convergence guarantees in terms of the $1$-Wasserstein distance. We provide an informal version of our Theorem 4.1 below.
@@
Consider a data domain $\Omega\subset\mathbb{R}^{d}$ with $\ell_{2}$ diameter $D$. For any dataset $S\in\Omega^{n}$, there exists a parameter setting such that PE (Algorithm 2) is $(\varepsilon,\delta)$-DP and it outputs a synthetic dataset $S^{\prime}$ satisfying
@@
As a corollary, we find sufficient conditions on the APIs used by PE to understand its convergence in more general settings, such as Banach spaces. We also identify strong connections between PE and the Private Signed Measure Mechanism (PSMM), an algorithm for DP synthetic data generation under pure DP [24] that has theoretical guarantees. Hence, our work bridges the theory and practice of PE in both directions: we provide theory for a practical version of PE, and show how PE naturally arises as a practical version of PSMM.
@@
Our work significantly improves the theory of PE presented in [31], providing a more realistic theoretically tractable version of PE and eliminating unrealistic assumptions in the convergence analysis. More concretely, we make the following contributions.
  • We prove a lower bound (Lemma 3.1) that establishes that without the multiplicity assumption on the data, the convergence proof of PE provided in [31] only works for undesirable privacy parameters under pure DP. This indicates that a new analysis for PE is needed.
  • We propose a new theoretically tractable variant of PE, Algorithm 2. Under this model, we formally prove worst-case convergence rates with respect to the $1$-Wasserstein distance as $|S|\to\infty$; see Theorem 4.1. We identify sufficient conditions for convergence of PE in more general settings such as Banach spaces (Appendix F).
  • We draw connections between PE and the Private Signed Measure Mechanism (PSMM), an algorithm for DP synthetic data generation [24]. We exploit this connection by using tools from the analysis of PSMM to prove the convergence of PE. Finally, we also show how PE arises naturally when trying to make PSMM ‘practical’. See Section 5 for details.
@@ 2 Preliminaries
Definition 2.1 (Differential Privacy [16]).
@@
Definition 2.2 ($1$-Wasserstein distance [40]).
@@
[…]
@@
In the past, there have been different theoretical formulations of the DP Synthetic Data problem (see Section A). In most of them, a fixed set of queries $\mathcal{Q}$ is used to measure the accuracy of the synthetic data: the answer to a query $q\in\mathcal{Q}$ should be similar when querying the sensitive dataset $S$ and the synthetic dataset $S^{\prime}$. In other words, $\max_{q\in\mathcal{Q}}|q(S)-q(S^{\prime})|$ should be small. By Kantorovich duality [40], we can alternatively define the $1$-Wasserstein distance as
@@
[…]
@@ 3 Background: Private Evolution (PE) and Prior Theory
Lin et al. present the only prior convergence analysis of PE [31]. Their analysis applies to a variant of PE (that we call theoretical PE), which is outlined in Algorithm 1. Below, we first describe Algorithm 1 and explain how it differs from how PE is used in practice (Algorithm 3, which we call practical PE). Then, we explain the limitations of prior analysis.
Algorithm 1 (Theoretical) Private Evolution; steps in blue differ from practical PE [31]
@@
PE (both the practical variant, and the theoretical variant in Algorithm 1) generates a DP synthetic dataset by iteratively refining an initial random dataset using a DP nearest-neighbor histogram (Algorithm 4). To do so, it requires access to two APIs that are independent of the sensitive dataset $S$, $\operatorname{Random\_API}$ and $\operatorname{Variation\_API}$ [^1] In practice, generative models trained on public data and simulators can be used as APIs..
@@
    where $\alpha>0$ is a small constant and $N^{k,l}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{l}^{2}I_{d})$ with $\sigma_{l}=\frac{\alpha 2^{l-1}}{\sqrt{\pi}[(\sqrt{d}+\log(2))^{2}+\log(2)]}$. Recall that for a dataset $S_{t}$, $\operatorname{Variation\_API}(S_{t})=\cup_{z\in S_{t}}\operatorname{Variation\_API}(z)$. That is, the algorithm generates $O(|S_{t}|\log(\text{diam}(\Omega)/\alpha))$ variations by adding independent, spherical Gaussian noise to each sample $z$, then projecting the sum back onto $\Omega$. Then, it returns a set containing the original $z$ and all generated variations.
@@
Practical PE and Algorithm 1 differ in significant ways, which are highlighted in blue in Algorithm 1. Most notably, Algorithm 1 creates the next synthetic dataset $S_{t}$ deterministically by adding variations to the next dataset proportionally to the entries of the DP histogram. In practical PE, the next synthetic dataset is instead constructed by sampling with replacement from a distribution defined by the histogram. This seemingly small difference is important for their proof technique, discussed next. More differences between the two variants of PE are discussed in Appendix B.
Limitations of the utility analysis of Algorithm 1 in [31].
The theoretical analysis in [31] relies on first showing that noiseless PE (Algorithm 1 with $\sigma=0$) converges, then arguing that when each data point from the sensitive dataset $S$ is repeated $B$ times, with high probability the noisy version behaves like noiseless PE for large $B$.
In the noiseless case the elements of $S$ and $S_{0}$ are matched: assign $S_{0}[i]$ to $S[i]$. Then iteratively, for each $i$, $S_{t}[i]$ is chosen as the element from $\operatorname{Variation\_API}(S_{t-1})$ that is closest to $S[i]$, which is likely to be closer to $S[i]$ than $S_{t-1}[i]$. After enough steps, $\rho(S[i],S_{T}[i])\leq\eta$ for all $i\in[n]$; we say the datasets are $\eta$-close, written as $S=_{\eta}S_{T}$. To handle noise, [31] assumes every real data point is repeated $B$ times. This boosts the signal in the noisy step (Line 7), allowing the algorithm to behave like the noiseless version. In other words, although the proof assume each data point is repeated $B$ times, the noise scale provides a DP guarantee for neighboring datasets that differ in a single sample. This setup is unrealistic and sidesteps the core challenge of differentially private learning. More details in Appendix B.
We provide a formal lower bound giving evidence that the proof technique in [31] is fundamentally limited, in the sense that their unrealistic assumption is necessary for their proof technique to work. More concretely, in Lemma 3.1 we show that if we remove the multiplicity assumption, then any $\varepsilon$-DP algorithm can only output a DP synthetic dataset $S_{T}$ that is $\eta$-close to $S$ when $\varepsilon=\Omega(d\log(1/\eta))$, or equivalently, $\eta=\Omega(e^{-\varepsilon/d})$.
@@
Consider a metric (data) space $(\Omega,\rho)$ and a fixed dataset $S\in\Omega^{n}$. Let $\eta>0$ and $\mathcal{A}:\Omega^{n}\to\Omega^{n}$ be such that $\mathbb{P}_{\mathcal{A}}\big{[}S=_{\eta}\mathcal{A}(S)\big{]}\geq 1-\tau$. Suppose $\mathcal{A}$ is $\varepsilon$-DP. Then, denoting $\mathcal{M}(\Omega,\rho;2\eta)$ the $2\eta$ packing number of $\Omega$ w.r.t $\rho$, we must have
$\varepsilon\geq\log((\mathcal{M}(\Omega,\rho;2\eta)-1)(1-\tau)/\tau).$
@@
The packing number $\mathcal{M}(\Omega,\rho;2\eta)$ is of the order $(1/\eta)^{d}$ when $\Omega\subset\mathbb{R}^{d}$ and $\operatorname{diam}(\Omega)\leq 2$ (see Lemma $5.5$ and Example $5.8$ in [41]).
@@ 4 Convergence of Private Evolution
We have established several limitations of prior utility analysis of PE: the algorithmic variant of PE analyzed in [31] differs from what is done in practice, and the analysis itself is unlikely to extend beyond very weak privacy regimes, as evidenced by Lemma 3.1. In this section, we introduce a new theoretically tractable version of PE along with its utility convergence guarantees. The algorithm we analyze more closely reflects how PE works in practice and is amenable to worst-case utility analysis.

### v2

@@ Abstract
Private Evolution (PE) is a promising training-free method for differentially private (DP) synthetic data generation. While it achieves strong performance in some domains (e.g., images and text), its behavior in others (e.g., tabular data) is less consistent. To date, the only theoretical analysis of the convergence of PE depends on unrealistic assumptions about both the algorithm’s behavior and the structure of the sensitive dataset. In this work, we develop a new theoretical framework to understand PE’s practical behavior and identify sufficient conditions for its convergence. For $d$-dimensional sensitive datasets with $n$ data points from a convex and compact domain, we prove that under the right hyperparameter settings and given access to the Gaussian variation API proposed in PE23, PE produces an $(\varepsilon,\delta)$-DP synthetic dataset with expected 1-Wasserstein distance $\tilde{O}(d(n\varepsilon)^{-1/d})$ from the original; this establishes worst-case convergence of the algorithm as $n\to\infty$. Our analysis extends to general Banach spaces as well. We also connect PE to the Private Signed Measure Mechanism, a method for DP synthetic data generation that has thus far not seen much practical adoption. We demonstrate the practical relevance of our theoretical findings in experiments.
@@ 1 Introduction
Many modern machine learning applications rely on user data, making data privacy protection a central concern. In this regard, differential privacy (DP) dwork2006:calibrating; dwork2014DPfoundations has become a de facto standard for safeguarding sensitive information of individuals. Given a private dataset, many problems—including regression sheffet2019DPlinearregression; chaudhuri2008DPlogisticregression, deep learning abadi2016DPdeeplearning, and stochastic optimization bassily2021DPSO; arora2023DPnonconvex—can be performed in a DP manner. While it is possible to design a different DP algorithm for each specific task, a reasonable alternative is to generate a DP synthetic dataset that preserves many of the statistical properties of the sensitive dataset. By the post-processing property of DP dwork2014DPfoundations, the DP synthetic data can then be fed into existing non-DP algorithms without incurring additional privacy loss; this avoids modifying existing ML pipelines and enables data sharing with third parties (e.g. for reproducibility).
Recently, PE23 introduced Private Evolution (PE), a promising new framework for DP synthetic data generation that relies on public, pretrained data generators pe2; pe3; hou24PE_federated; swanberg2025PEtabulardata; zou2025PEmultipleLLMs; hou2025PEandPO. PE is currently competitive with—and sometimes improves on—state-of-the-art models in terms of Fréchet inception distance (FID) and downstream task performance in settings such as images and text PE23; pe2; pe3; hou24PE_federated. In addition, PE is training-free, whereas current state-of-the-art approaches typically train (or finetune) a generative model on the sensitive dataset using DP-SGD xie2018DPgan; long2021gpate; dockhorn2022dpdifussionmodels; cao2021DPGMwithsinkhorn. However, in some settings, including tabular data swanberg2025PEtabulardata, and image data with mismatched distributions gong2025dpimagebench, PE has achieved limited success. To better understand when PE works, it is crucial to improve our theoretical understanding of the algorithm.
@@
In PE23, the authors study a theoretically tractable variant of PE, analyzing its convergence in terms of the Wasserstein distance, with the aim of understanding its behavior and justifying its empirical success. However, this theoretically tractable version of PE differs in many significant ways from the algorithm that is used in experiments; further, the convergence analysis in PE23 makes the unrealistic assumption that every point in $S$ is repeated many times. We start by showing that the proof technique in PE23 is inherently limited by this multiplicity assumption; if we remove the assumption, then PE can only be $\varepsilon$-DP with a parameter $\varepsilon$ that scales as $d\log(1/\eta)$, where $d$ is the ambient dimension of the data and $\eta$ is the final accuracy of the algorithm in terms of the Wasserstein distance. This is impractical in even moderate dimensions.
In this paper, we address the limitations in prior analysis of PE by providing a convergence analysis that does not require strong assumptions about the nature of the dataset and more closely matches the PE algorithm used in practice. We formally prove worst-case [^1] Note that convergence is proven in the worst-case over problem instances for a specific set of hyperparameters and APIs. convergence guarantees in terms of the $1$-Wasserstein distance. We provide an informal version of our Theorem 4.1 below.
@@
Consider a data domain $\Omega\subset\mathbb{R}^{d}$ with $\ell_{2}$ diameter $D$. For any dataset $S\in\Omega^{n}$ and $0<\varepsilon,\delta<1$, there exist APIs and a parameter setting such that PE (Algorithm 2) is $(\varepsilon,\delta)$-DP and it outputs a synthetic dataset $S^{\prime}$ satisfying
@@
As a corollary, we find sufficient conditions on the APIs used by PE to understand its convergence in more general settings, such as Banach spaces. We also identify strong connections between PE and the Private Signed Measure Mechanism (PSMM), an algorithm for DP synthetic data generation under pure DP Vershynin23DPsyntheticdata that has theoretical guarantees. Hence, our work bridges the theory and practice of PE in both directions: we provide theory for a practical version of PE, and show how PE naturally arises as a practical version of PSMM.
@@
Our work significantly improves the theory of PE presented in PE23, providing a more realistic theoretically tractable version of PE and eliminating unrealistic assumptions in the convergence analysis. More concretely, we make the following contributions.
  • We prove a lower bound (Lemma 3.1) that establishes that without the multiplicity assumption on the data, the convergence proof of PE provided in PE23 only works for undesirable privacy parameters under pure DP. This indicates that a new analysis for PE is needed.
  • We propose a new theoretically tractable variant of PE, Algorithm 2. Under this model, we formally prove worst-case convergence rates with respect to the $1$-Wasserstein distance as $|S|\to\infty$; see Theorem 4.1. We identify sufficient conditions for convergence of PE in more general settings such as Banach spaces (Appendix G).
  • We draw connections between PE and the Private Signed Measure Mechanism (PSMM), an algorithm for DP synthetic data generation Vershynin23DPsyntheticdata. We exploit this connection by using tools from the analysis of PSMM to prove the convergence of PE. Finally, we also show how PE arises naturally when trying to make PSMM ‘practical’. See Section 5 for details.
@@ 2 Preliminaries
Definition 2.1 (Differential Privacy dwork2006:calibrating).
@@
Definition 2.2 ($1$-Wasserstein distance villani2008optimal).
@@
$W_{1}(\mu,\nu)=\inf_{\gamma\in\Gamma(\mu,\nu)}\int_{\Omega\times\Omega}\rho(x,y)d\gamma(x,y),$
@@
In the past, there have been different theoretical formulations of the DP Synthetic Data problem (see Section A). In most of them, a fixed set of queries $\mathcal{Q}$ is used to measure the accuracy of the synthetic data: the answer to a query $q\in\mathcal{Q}$ should be similar when querying the sensitive dataset $S$ and the synthetic dataset $S^{\prime}$. In other words, $\max_{q\in\mathcal{Q}}|q(S)-q(S^{\prime})|$ should be small. By Kantorovich duality villani2008optimal, we can alternatively define the $1$-Wasserstein distance as
@@
Samples lie in the metric space $(\Omega,\rho)$. The diameter of $\Omega$ is denoted by $\operatorname{diam}(\Omega)=\sup_{z_{1},z_{2}\in\Omega}\rho(z_{1},z_{2})$. $\mathcal{P}(\Omega)$ is the space of probability measures supported on $\Omega$. A dataset is a set of elements from $\Omega$. For a dataset $S$, $|S|$ represents its cardinality, $S[i]$ its $i$-th element and $\mu_{S}$ the empirical distribution $\frac{1}{|S|}\sum_{z\in S}\delta_{z}$, where $\delta_{z}$ is the dirac delta mass at $z$. The bounded Lipschitz distance between the signed measures $\mu,\nu$ supported in $\Omega$ is $D_{BL}(\mu,\nu):=\sup_{f\in{\cal F}_{BL}}\int fdu-\int fdv$. For probability measures $\mu,\nu$, $D_{BL}(\mu,\nu)=W_{1}(\mu,\nu)$. $\Delta_{d}$ is the standard probability simplex in $\mathbb{R}^{d}$. $v[i]$ indicates the $i$-th coordinate of a vector $v\in\mathbb{R}^{d}$. $\texttt{nint}(x)$ denotes the nearest integer to $x$.
@@ 3 Background: Private Evolution (PE) and Prior Theory
Lin et al. present the only prior convergence analysis of PE PE23. Their analysis applies to a variant of PE (that we call theoretical PE), which is outlined in Algorithm 1. Below, we first describe Algorithm 1 and explain how it differs from how PE is used in practice (Algorithm 3, which we call practical PE). Then, we explain the limitations of prior analysis.
Algorithm 1 (Theoretical) Private Evolution; steps in blue differ from practical PE PE23
@@
PE (both the practical variant, and the theoretical variant in Algorithm 1) generates a DP synthetic dataset by iteratively refining an initial random dataset using a DP nearest-neighbor histogram (Algorithm 4). To do so, it requires access to two APIs that are independent of the sensitive dataset $S$, $\operatorname{Random\_API}$ and $\operatorname{Variation\_API}$ [^2] In practice, generative models trained on public data and simulators can be used as APIs.. For our theory, we use the same theoretical API models proposed in PE23, which we detail below.
@@
    where $\alpha>0$ is a small constant and $N^{k,l}\overset{iid}{\sim}\mathcal{N}(0,\sigma_{l}^{2}I_{d})$ with $\sigma_{l}=\frac{\alpha 2^{l-1}}{\sqrt{\pi}[(\sqrt{d}+\log(2))^{2}+\log(2)]}$. Note that noise is added at different scales. This is key for the theory to work, because it allows to ‘explore’ the data domain more effectively. Recall that for a dataset $S_{t}$, $\operatorname{Variation\_API}(S_{t})=\cup_{z\in S_{t}}\operatorname{Variation\_API}(z)$. That is, the API generates $O(|S_{t}|\log(\text{diam}(\Omega)/\alpha))$ variations by adding independent, spherical Gaussian noises with increasing variances to each sample $z$.
@@
Practical PE and Algorithm 1 differ in significant ways, which are highlighted in blue in Algorithm 1. Most notably, Algorithm 1 creates the next synthetic dataset $S_{t}$ deterministically by adding variations to the next dataset proportionally to the entries of the DP histogram. In practical PE, the next synthetic dataset is instead constructed by sampling with replacement from a distribution defined by the histogram. This seemingly small difference is important for their proof technique, discussed next. More differences between the two variants of PE are discussed in Appendix C.
Limitations of the utility analysis of Algorithm 1 in PE23.
The theoretical analysis in PE23 relies on first showing that noiseless PE (Algorithm 1 with $\sigma=0$) converges, then arguing that when each data point from the sensitive dataset $S$ is repeated $B$ times, with high probability the noisy version behaves like noiseless PE for large $B$.
In the noiseless case the elements of $S$ and $S_{0}$ are matched: assign $S_{0}[i]$ to $S[i]$. Then iteratively, for each $i$, $S_{t}[i]$ is chosen as the element from $\operatorname{Variation\_API}(S_{t-1})$ that is closest to $S[i]$, which is likely to be closer to $S[i]$ than $S_{t-1}[i]$. After enough steps, $\rho(S[i],S_{T}[i])\leq\eta$ for all $i\in[n]$; we say the datasets are $\eta$-close, written as $S=_{\eta}S_{T}$. To handle noise, PE23 assumes every real data point is repeated $B$ times. This boosts the signal in the noisy step (Line 7), allowing the algorithm to behave like the noiseless version. In other words, although the proof assume each data point is repeated $B$ times, the noise scale provides a DP guarantee for neighboring datasets that differ in a single sample. This setup is unrealistic and sidesteps the core challenge of differentially private learning. More details in Appendix C.
We provide a formal lower bound giving evidence that the proof technique in PE23 is fundamentally limited, in the sense that their unrealistic assumption is necessary for their proof technique to work. More concretely, in Lemma 3.1 we show that if we remove the multiplicity assumption, then any $\varepsilon$-DP algorithm can only output a DP synthetic dataset $S_{T}$ that is $\eta$-close to $S$ when $\varepsilon=\Omega(d\log(1/\eta))$, or equivalently, $\eta=\Omega(e^{-\varepsilon/d})$. Proof in Appendix H.
@@
Consider a metric (data) space $(\Omega,\rho)$ and a fixed dataset $S\in\Omega^{n}$. Let $\eta>0$ and $\mathcal{A}:\Omega^{n}\to\Omega^{n}$ be such that $\mathbb{P}_{\mathcal{A}}\big[S=_{\eta}\mathcal{A}(S)\big]\geq 1-\tau$ for some $\tau\in(0,1/4)$. Suppose $\mathcal{A}$ is $(\varepsilon,\delta)$-DP for some $\delta\in(0,1-3\tau)$. Then, denoting $\mathcal{M}(\Omega,\rho;2\eta)$ the $2\eta$ packing number of $\Omega$ w.r.t $\rho$, we must have
$\varepsilon\geq\log(\mathcal{M}(\Omega,\rho;2\eta)).$
@@
The packing number $\mathcal{M}(\Omega,\rho;2\eta)$ is of the order $(1/\eta)^{d}$ when $\Omega\subset\mathbb{R}^{d}$ and $\operatorname{diam}(\Omega)\leq 2$ (see Lemma $5.5$ and Example $5.8$ in wainwright2019HDS).
@@ 4 Convergence of Private Evolution
We have established several limitations of prior utility analysis of PE: the algorithmic variant of PE analyzed in PE23 differs from what is done in practice, and the analysis itself is unlikely to extend beyond very weak privacy regimes, as evidenced by Lemma 3.1. In this section, we introduce a new theoretically tractable version of PE along with its utility convergence guarantees. The algorithm we analyze more closely reflects how PE works in practice and is amenable to worst-case utility analysis.